# Mutation-acyclic quivers are totally proper

### v1

@@ 6. Applications
We next illustrate a way of combining Corollary 6.6 with another mutation invariant [20], the multiset of GCDs of all the integers in the columns (or rows) of the $B$-matrix. We construct a COQ as in Example 6.12 with $a=b=2$ and $c=4$. Thus $M_{Q}=12\geq 3$, and we cannot use Corollary 6.6 to conclude that $Q$ is not mutation-acyclic. However, note that the GCD of the entries of $B_{Q}$ is $2$. Thus if $Q$ is mutation-acyclic, it must be mutation equivalent to a connected acyclic quiver with the same GCD and $M_{Q}=12$. There are two such quivers, both trees with $2$ arrows between each pair of neighboring vertecies. The $B$-matrices of these acyclic quivers have determinant 00 and $16$, but $\det(B_{Q})=144$. Thus $Q$ is not mutation-acyclic.
@@
Fix a totally proper COQ $Q$ and a proper vertex $v_{j}$. Let $U_{Q}$ be a unipotent companion with linear order $<$ so that $v_{i}<v_{j}$ (resp., $v_{j}<v_{i}$) for all $v_{i}\in{\operatorname{In}}(v_{j})$ (resp., ${\operatorname{Out}}(v_{j})$). Then $M_{Q}(v_{j},\epsilon)(U_{Q}+U_{Q}^{T})M_{Q}(v_{j},\epsilon)$ is an admissible quasi-Cartan companion of $\mu_{v_{j}}(Q)$ when $\epsilon=-1$ (resp., $1$).

### v2

@@ 6. Applications
We next illustrate a way of combining Corollary 6.7 with another mutation invariant [23], the multiset of GCDs of all the integers in the columns (or rows) of the $B$-matrix. We construct a COQ as in Example 6.14 with $a=b=2$ and $c=4$. Thus $M_{Q}=12\geq 3$, and we cannot use Corollary 6.6 to conclude that $Q$ is not mutation-acyclic. However, note that the GCD of the entries of $B_{Q}$ is $2$. Thus if $Q$ is mutation-acyclic, it must be mutation equivalent to a connected acyclic quiver with the same GCD and $M_{Q}=12$. There are two such quivers, both trees with $2$ arrows between each pair of neighboring vertecies. The $B$-matrices of these acyclic quivers have determinant 0 and $16$, but $\det(B_{Q})=144$. Thus $Q$ is not mutation-acyclic.
It is natural to ask when the Alexander polynomial is unimodal (say up to the signs of the coefficients). If we make no assumptions on the quiver, then the general answer seems complicated (note that the Alexander polynomial of the quivers in Example 6.13 are not). We give two examples from acyclic quivers.
Example 6.16.
Let $Q$ be the COQ shown below:
A COQ with cyclic ordering (v1,v2,v3,v4) and 2 arrows each between v1 and v2, v2 and v3, and v1 and v3, as well as 4 arrows from v3 to v4.
Then $\Delta_{Q}(t)=t^{4}+32t^{3}-2t^{2}+32t+1$ is not unimodal.
Example 6.17.
Consider the COQ $Q$ below:
A COQ with cyclic ordering (v1,v2,v3,v4). There are a1 arrows from v1 to v3, a2 from v2 to v3, a3 from v1 to v4, and a4 from v2 to v4.
We compute the Alexander polynomial
$\Delta_{Q}(t)=$ $t^{4}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}-4)t^{3}$
$+(a_{2}^{2}a_{3}^{2}-2a_{1}a_{2}a_{3}a_{4}+a_{1}^{2}a_{4}^{2}-2a_{1}^{2}-2a_{2}^{2}-2a_{3}^{2}-2a_{4}^{2}+6)t^{2}$
$+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}-4)t+1$
$=$ $(t-1)^{4}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2})t(t-1)^{2}+(a_{2}a_{3}-a_{1}a_{4})^{2}t^{2}.$
A polynomial is $\gamma$-positive if it can be expressed as $\sum_{i}\gamma_{i}t^{i}(t+1)^{n-2i}$ for some nonnegative coefficients $\gamma_{i}$ [1]. Our second expression above is almost of this form, except that we have (t1)(t-1) instead of $(t+1)$. It is not hard to show that $\Delta_{Q}(t)$ will be unimodal (indeed, $\gamma$-positive) when $64<4(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2})<(a_{2}a_{3}-a_{1}a_{4})^{2}$.
@@
Fix a totally proper COQ $Q$ and a proper vertex $v_{j}$. Let $U_{Q}$ be a unipotent companion with linear order << so that $v_{i}<v_{j}$ (resp., $v_{j}<v_{i}$) for all $v_{i}\in{\operatorname{In}}(v_{j})$ (resp., ${\operatorname{Out}}(v_{j})$). Then $M_{Q}(v_{j},\epsilon)(U_{Q}+U_{Q}^{T})M_{Q}(v_{j},\epsilon)$ is an admissible quasi-Cartan companion of $\mu_{v_{j}}(Q)$ when $\epsilon=-1$ (resp., $1$).